OpenStudy (anonymous):

I have a small problem with the Recitaion video for elimination with matrices. With the second pivot (2), she uses 1/2 as a multiplier. However, in my own solution I multiply by -1/2. Here is my logic: If we are to change the -1 in row 3 to 0, we need to subtract -1 from it. Therefore multiplying the pivot by -1/2 gets us -1 and therefore 0, not 1/2. The solution in the video is -1 - (1/2*2) = -2 not 0 Am I completely wrong or is there a mistake on the video?

OpenStudy (anonymous):

Oh, I should get more sleep, didn't realize that row2 = row2 - (-1/2 row1) is the dame as row2 = row2 + 1/2row1